40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog, Page 1 of 9  Not logged in ELOG logo
ID Date Author Type Category Subject
  411   Wed Jul 7 14:21:50 2021 StephenSupplyGeneralOMC Unit 4 Build Machined Parts

More explicit insights into the inventory for the Unit 4 build. Image of inventory included below.

Machined Parts:

Cable Components:

  • Hughes Circuits made us Kapton flex circuits. These have not been processed in any way.
  • Rich had supplied a spool of Gore 4-conductor in-vacuum wire (see below image). I returned the sppol for Rich but it is living in Downs and available for use.
  • PEEK cable ties were damaged during bake, and will be replaced by SYS inventory.

Retrofit/Repair Capabilities:

  • Aluminum reinforcement brackets D1600316
  • Glass reinforcement brackets (Edmund Optics 45-072 and 45-071)

ref: E1900034 and other associated documents.

Attachment 1: IMG_9238.JPG
IMG_9238.JPG
Attachment 2: IMG_9236.JPG
IMG_9236.JPG
  410   Sun May 30 15:32:56 2021 KojiGeneralGeneralDCPD AF capacitance measirement

Measurement result:

The capacitance at no bias was 460~500pF. This goes down to below 300pF at 1.0~1.5V reverse bias. At maximum +15V, the capacitance goes down to 200~220pF.

On this opportunity, the capacitances of a couple of Excelitas C30665 photodiodes were measured. In Attachment 2, the result was compared with one of the results from the high QE PDs. In general the capacitance of C30665 is lower than the one from the high QE PDs.

Attachment 1: highQEPD_capacitance.pdf
highQEPD_capacitance.pdf
Attachment 2: C30665_capacitance.pdf
C30665_capacitance.pdf
  409   Sun May 30 15:17:16 2021 KojiGeneralGeneralDCPD AF capacitance measirement

Attachment 1: System diagram. The reverse bias voltage is controlled by DS335. This can produce a voltage offset up to 10V. A G=+2 opamp circuit was inserted so that a bias of up to +15V can be produced. The capacitances of the photodiodes were measured with SR720 LCR meter with a probe. DS335 and SR720 were controlled from PC/Mac via serial connections.

Attachment 2: Overview

Attachment 3: How was the probe attached to the photodiode under the test

Attachment 4: The bias circuitry and the power supply

Attachment 5: G=+2 amp

Attachment 1: PD_cap_meas.pdf
PD_cap_meas.pdf
Attachment 2: 20210529013015_IMG_0577.jpeg
20210529013015_IMG_0577.jpeg
Attachment 3: 20210529013114_IMG_0580_2.jpeg
20210529013114_IMG_0580_2.jpeg
Attachment 4: 20210529013200_IMG_0584.jpeg
20210529013200_IMG_0584.jpeg
Attachment 5: 20210529013229_IMG_0586.jpeg
20210529013229_IMG_0586.jpeg
  408   Thu May 20 17:03:50 2021 KojiGeneralGeneralSRS LCR meter SRS720 borrowed from Downs

Item loan: SRS LCR meter SRS720 borrowed from Downs. The unit is at the 40m right now for testing with an excelitas PD. Once it is done, the setup will be moved to the OMC lab for testing the high QE PDs

Attachment 1: P_20210520_154841.jpg
P_20210520_154841.jpg
  407   Fri Feb 5 07:40:37 2021 StephenSupplyGeneralOMC Unit 4 Build Machined Parts

OMC Unit 4 Build Machined Parts are currently located in Stephen's office. See image of large blue box from office, below.

Loaned item D1100855-V1-00-OMC08Q004 to Don Griffith for work in semi-clean HDS assy.

This includes mass mounting brackets, cable brackets, balance masses, etc. For full inventory, refer to ICS load Bake-9527 (mixed polymers) and Bake-9495 (mixed metals).

Inventory includes all items except cables. Plasma sprayed components with slight chipping were deemed acceptable for Unit 4 use. Cable components (including flex circuit) are ready to advance to fabrication, with a bit more planning and ID of appropriate wiring.

Attachment 1: IMG_8117.JPG
IMG_8117.JPG
  406   Tue Nov 24 12:27:18 2020 KojiElectronicsCharacterizationThe dark noise of the Q3000 QPDs

The amplifier BW was 400kHz at the gain of 1e7 V/A. And the max BW is 500kHz even at a lower gain. I have to setup something special to see the RF band dark noise.
With this situation, I stated "the RF dark noise should be characterized by the actual WFS head circuit." in the 40m ELOG.

  405   Tue Nov 24 10:45:07 2020 gautamElectronicsCharacterizationThe dark noise of the Q3000 QPDs

I see that these measurements are done out to 100 kHz - I guess there is no reason to suspect anything at 55 MHz which is where this QPD will be reading out photocurrent given the low frequency behavior looks fine? The broad feature at ~80 kHz is the usual SR785 feature I guess, IIRC it's got to do with the display scanning rate.

Quote:

The measured floor level of the dark current was below the shot noise level for the DC current of 0.1mA (i.e. 6pA/rtHz).

  404   Mon Nov 23 23:17:19 2020 KojiElectronicsCharacterizationThe dark noise of the Q3000 QPDs

The dark noise levels of the four Q3000 QPDs were measured with FEMTO DLPCA200 low noise transimpedance amp.

The measurement has been done in the audio frequency band. The amp gain was 10^7 V/A. The reverse bias was set to be 5V and the DC output of the amplifier was ~40mV which corresponds to the dark current of 4nA. It is consistent with the dark current measurement.

The measured floor level of the dark current was below the shot noise level for the DC current of 0.1mA (i.e. 6pA/rtHz).
No anomalous behavior was found with the QPDs.

Note that there is a difference in the level of the power line noise between the QPDs. The large part of the line noises was due to the noise coupling from a soldering iron right next to the measurement setup, although the switch of the iron was off. I've noticed this noise during the measurement sets for QPD #83. Then the iron was disconnected from the AC tap.
 

Attachment 1: Q3000_dark_noise_81.pdf
Q3000_dark_noise_81.pdf
Attachment 2: Q3000_dark_noise_82.pdf
Q3000_dark_noise_82.pdf
Attachment 3: Q3000_dark_noise_83.pdf
Q3000_dark_noise_83.pdf
Attachment 4: Q3000_dark_noise_84.pdf
Q3000_dark_noise_84.pdf
  403   Sun Nov 22 13:49:12 2020 KojiElectronicsCharacterizationImpedance Measurement for InGaAs QPDs

To know any anomaly to the junction capacitance of the QPD segments, the RF impedances were tested with a hand-made impedance measurement.
All segments look almost identical in terms of capacitance.

Measurement setup:
The impedance of a device can be measured, for example, from the complex reflection coefficient (S11). To measure the reflection, a bidirectional coupler was brought from the 40m. Attachments 1 and 2 shows the connection. The quantity A/R shows S11. The network analyzer can convert a raw transfer function to an impedance in Ohm.

Calibration and Measurement limit:
The network analyzer was calibrated with 1) a piece of wire to short the clips 2) 50ohm resistor 3) open clips. Then the setup was tested with these three conditions (again). Attachment 3 shows the result. Because of the impedance variation of the system (mainly from the Pomona clip, I guess), there looks the systematic measurement error of ~1pF or ~25nH. Above 100MHz, the effect of the stray impedance is large such that the measurement is not reliable.

The setup was tested with a 10pF ceramic capacitor and this indicated it is accurate at this level. The setup is sufficient for measuring the diode junction capacitance of 300~500pF.

Impedance of the QPD segments:

Then the impedances of the QPD segments were measured (Attachment 4). The segments showed the identical capacitance of 300~400pF level, except for the variation of the stray inductance at high freq, which we can ignore. Note that there is no bias voltage applied and the nominal capacitance in the datasheet is 225pF at 5V reverse bias. So I can conclude that the QPDs are quite nominal in terms of the junction capacitance.

(Ed: 11/23/2020 The RF components were returned to the 40m)

Attachment 1: impedance_measurement.pdf
impedance_measurement.pdf
Attachment 2: P_20201121_183830.jpg
P_20201121_183830.jpg
Attachment 3: impedance_test.pdf
impedance_test.pdf
Attachment 4: Q3000_impedance_test.pdf
Q3000_impedance_test.pdf
  402   Sat Nov 21 13:58:30 2020 KojiElectronicsCharacterizationDark Current Measurement for InGaAs QPDs

Dark current measurement for InGaAs QPDs (OSI FCI-InGaAs-Q3000) has been done using Keithley 2450 and Frank's diode test kit. Frank's setup uses various custom instruments which are no longer exist, therefore the kit was used only for switching between the segments.

The diodes were serialized as 81, 82, 83, 84, continuing the numbering for the OMC QPDs. The numbers are engraved at the side and the back of the diode cans.

Overall, the QPDs nominally indicated the usual dark current level of <10nA.
SEG1 of #82 showed a lower voltage of reverse breakdown but this is not a critical level.
#83 showed variations between the segments compared to the uniform characteristics of #81 and #84.

Attachment 1: Q3000_dark_current.pdf
Q3000_dark_current.pdf
  401   Fri Nov 20 18:51:23 2020 KojiGeneralGeneralInstrument loan

FEMTO DLPCA200 low noise preamp (brand new)

Keithley Source Meter 2450 (brand new) => Returned 11/23/2020

were brought to the OMC lab for temporary use.

https://nodus.ligo.caltech.edu:8081/QIL/2522

  400   Mon Nov 9 22:06:18 2020 KojiMechanicsGeneral5th OMC Transport Fixture

I helped to complete the 5th OMC Transport Fixture. It was built at the 40m clean room and brought to the OMC lab. The fixture hardware (~screws) were also brought there.

Attachment 1: IMG_0211.jpg
IMG_0211.jpg
Attachment 2: IMG_0221.jpg
IMG_0221.jpg
  399   Fri Nov 6 18:38:00 2020 KojiGeneralGeneralPowermeter lent from OMC Lab to 2um ECDL

Thorlabs' powermeter controler + S401C head was lent from OMC Lab.

https://nodus.ligo.caltech.edu:8081/SUS_Lab/1856

  398   Fri Oct 23 19:09:54 2020 KojiGeneralGeneralParticle counter transfered to Radhika

See this entry: https://nodus.ligo.caltech.edu:8081/40m/15642

  397   Fri Oct 16 00:53:29 2020 KojiGeneralGeneralTFT Monitor mounting

The image flipping of the display unit was fixed. The vendor told me how to fix it.

- Open the chassis by the four screws at the side.
- Look at the pass-through PCB board between the mother and display boards.
- Disconnect the flat flex cables from the pass-through PCB (both sides) and reconnect them (i.e. reseat the cables)

That's it and it actually fixed the image flipping issue.

 

  396   Fri Oct 9 01:01:01 2020 KojiGeneralGeneralTFT Monitor mounting

To spare some room on the optical table, I wanted to mount the two TFT monitor units on the HEPA enclosure frame.
I found some Bosch Rexroth parts (# 3842539840) in the lab, so the bracket was taken for the mount. This swivel head works very well. It's rigid and still the angle is adjustable.

https://www.boschrexroth.com/ics/cat/?cat=Assembly-Technology-Catalog&p=p834858

BTW, this TFT display (Triplett HDCM2) is also very nice. It has HDMI/VGA/Video/BNC inputs (wow perfect) and the LCD is Full-HD LED TFT.
https://www.triplett.com/products/cctv-security-camera-test-monitor-hd-1080p-led-display-hdcm2

https://www.newegg.com/p/0AF-0035-00016

https://www.bhphotovideo.com/c/product/1350407-REG/triplett_hdcm2_ultra_compact_7_hd_monitor.html

The only issue is that one unit (I have two) shows the image horizontally flipped. I believe that I used the unit with out this problem before, I'm asking the company how to fix this.

 

Attachment 1: 20201008214515_IMG_0152.jpg
20201008214515_IMG_0152.jpg
Attachment 2: 20201008214519_IMG_0153.jpg
20201008214519_IMG_0153.jpg
Attachment 3: 20201008214536_IMG_0154.jpg
20201008214536_IMG_0154.jpg
Attachment 4: 20201008220955_IMG_0155.jpg
20201008220955_IMG_0155.jpg
Attachment 5: 20201008221019_IMG_0156_2.jpg
20201008221019_IMG_0156_2.jpg
  395   Thu Oct 8 19:55:22 2020 KojiGeneralCharacterizationPower Measurement of Mephisto 800NE 1166A

The output of Mephisto 800NE (former TNI laser) was measured.
The output power was measured with Thorlabs sensors (S401C and S144C). The reference output record on the chassis says the output was 837mW at 2.1A injection.
They all showed some discrepancy. Thus we say that the max output of this laser is 1.03W at 2.1A injection based on the largest number I saw.

Attachment 1: Mephisto800NE_1166A.pdf
Mephisto800NE_1166A.pdf
  394   Mon Sep 28 16:13:08 2020 KojiGeneralGeneralOMC Beam Dump Production Cure Bake

According to the past backscatter test of the OMC (and the black glass beamdump: not V type but triangular type on a hexagonal-mount), the upper limit of the back reflection was 0.13ppm. https://nodus.ligo.caltech.edu:8081/OMC_Lab/209

I don't have a BRDF measurement. We can send a few black glass pieces to Josh.

  393   Mon Sep 28 16:03:13 2020 ranaGeneralGeneralOMC Beam Dump Production Cure Bake
are there any measurements of the BRDF of these things? I'm curious how much light is backscattered into the incoming beam and how much goes out into the world.

Maybe we can take some camera images of the cleaned ones or send 1-2 samples to Josh. No urgency, just curiosity.

I saw that ANU and also some labs in India use this kind of blue/green glass for beam dumps. I don't know much about it, but I am curious about its micro-roughness and how it compares to our usual black glass. For the BRDF, I think the roughnesss matters more for the blackness than the absorption.

  392   Mon Aug 10 15:53:46 2020 KojiGeneralGeneralLab status check

Check-in to the OMC lab to see the status. Nothing seemed changed. No bug. The HEPA is running normal. The particle level was 0.

Went into the HEPA enclosure and put a cover on the OMC. Because of the gluing template, the lid could not be close completely (that's expected and fine).

The IPA vector cloth bag was not dry yet but seemed expired (some smell). There is no stock left -> 5 bags to be ordered.

  391   Mon Aug 10 15:34:04 2020 KojiFacilityLoan / LendingGlue bake oven

Black and Decker Glue Baking Oven came back to the OMC lab on Aug 10, 2020, Georgia had lent the unit for the SAMS assembly/testing.

  390   Mon Aug 10 15:29:54 2020 KojiGeneralGeneralItem lending

The particle counter came back to the OMC lab on Aug 10, 2020

 

  389   Thu Feb 27 14:31:13 2020 KojiGeneralGeneralItem lending

Item lending as per Ian's request: Particle Counter from OMC Lab to QIL

 

Attachment 1: P_20200227_134755_vHDR_On.jpg
P_20200227_134755_vHDR_On.jpg
  388   Wed Dec 18 21:54:53 2019 KojiGeneralGeneralOMC Beam Dump Production Cure Bake

The beamdumps were taken out from the oven and packed in bags.

The bottom of the V are completely "wet" for 17 BDs among 20 (Attachment 1/2).

3 BDs showed insufficient glue or delamination although there is no sign of lack of rigidity. They were separated from the others in the pack.

Attachment 1: P_20191218_160650_vHDR_On.jpeg
P_20191218_160650_vHDR_On.jpeg
Attachment 2: P_20191218_160705_vHDR_On.jpeg
P_20191218_160705_vHDR_On.jpeg
Attachment 3: P_20191218_160733_003.jpeg
P_20191218_160733_003.jpeg
  387   Fri Dec 13 14:59:18 2019 StephenGeneralGeneralOMC Beam Dump Production Cure Bake

[Koji, Jordan, Stephen]

The beam dumps, bonded on Fri 06 Dec 2019, were placed in the newly tuned and configured small dirty ABO at the Bake Lab on Fri 13 Dec 2019.

Images are shared and references are linked below

Bonding log entry - https://nodus.ligo.caltech.edu:8081/OMC_Lab/386

Bake ticket - https://services.ligo-wa.caltech.edu/clean_and_bake/request/992/

OMC Beam Dump - https://dcc.ligo.org/LIGO-D1201285

Attachment 1: IMG_6080.JPG
IMG_6080.JPG
Attachment 2: IMG_6079.JPG
IMG_6079.JPG
  386   Fri Dec 6 00:55:25 2019 KojiOpticsGeneralBeamdump gluing

[Stephen, Koji]

20 glass beamdumps were bonded at the 40m cleanroom.


Attachment 1: We had 20 fused silica disks with a V-groove and 40 black glass pieces
Attachment 2: The black glass pieces had (usual) foggy features. It is well known to be very stubborn. We had to use IPA/acetone and wiping with pressure. Most of the feature was removed, but we could still see some. We decided to use the better side for the inner V surfaces.
Attachment 3: EP30-2 expiration date was 1/22/2020 👍. 7.66g of EP30-2 was poured and 0.38g of glass sphere was added. Total glue weight was 8.04g
Attachment 4: Glue test piece was baked at 200F in a toaster oven for ~12min. It had no stickiness. It was totally crisp. 👍👍👍
Attachment 5: Painted glue on the V-groove and put the glass pieces in. Then gave a dub of blue at the top and bottom of the V from the outside. In the end, we mostly had the glue went through the V part due to capillary action.
Attachment 6: The 20 BDs were stored in stainless vats. We looked at them for a while to confirm there is no drift and opening of the V part. Because the air bake oven was not available at the time, we decided to leave the assys there for the room temp curing, and then later bake them for the completion of the curing.
 

Attachment 1: 20191205114336_IMG_9171_1.jpeg
20191205114336_IMG_9171_1.jpeg
Attachment 2: 20191205114538_IMG_9173_1.jpeg
20191205114538_IMG_9173_1.jpeg
Attachment 3: 20191205161458_IMG_9175_1.jpeg
20191205161458_IMG_9175_1.jpeg
Attachment 4: 20191205163305_IMG_9183_1.jpeg
20191205163305_IMG_9183_1.jpeg
Attachment 5: 20191205172409_IMG_9187_1.jpeg
20191205172409_IMG_9187_1.jpeg
Attachment 6: 20191205172432_IMG_9188_1.jpeg
20191205172432_IMG_9188_1.jpeg
  385   Tue Oct 22 15:54:59 2019 KojiElectronicsLoan / LendingBorrowed LB1005 from Cryo Cav

From Cryo Cav setup

Borrowed LB1005 Servo box -> OMC

 

  384   Tue Oct 22 11:56:09 2019 StephenSupplyGeneralEpoxy Status update as of 22 October 2019

The following is the current status of the epoxies used in assembly of the OMC (excerpt from C1900052)

Re-purchasing efforts are underway and/or complete

Attachment 1: image_ep30-2_epoxy_kit_pcs_item_1582_location_downs_3303.JPG
image_ep30-2_epoxy_kit_pcs_item_1582_location_downs_3303.JPG
  383   Tue Oct 22 11:52:53 2019 StephenGeneralGeneralEpoxy Curing Timeline of OMC PZT Assy #9 and #10

This post captures the curing timeline followed by OMC PZT Assys #9 and #10.

Source file posted in case any updates or edits need to be made.

Attachment 1: omc_elog_383_Epoxy_Curing_Timeline_of_OMC_PZT_Assy_20191022.png
omc_elog_383_Epoxy_Curing_Timeline_of_OMC_PZT_Assy_20191022.png
Attachment 2: omc_elog_383_Epoxy_Curing_Timeline_of_OMC_PZT_Assy.pptx
  382   Tue Oct 22 10:25:01 2019 StephenGeneralGeneralOMC PZT Assy #9 and #10 Production Cure Bake

OMC PZT Assy Production Cure Bake (ref. OMC elog 381) for PZT Assy #9 and #10 started 27 September 2019 and completed 28 September 2019. Captured in the below figure (purple trace). Raw data has been posted as an attachment as well.

We have monitored the temperature in two ways:

1) Datalogger thermocouple data (purple trace).
2) Checking in on temperature of datalogger thermocouple (lavender circles) and drive thermocouple (lavender diamonds), only during initial ramp up.

Comments on bake:

  • No changes were made to the tuning or instrumentation of the oven between the successful qualifying bake obtained on 26 September (ref. OMC elog 380). However, the profile seems to have been more similar to prior qualifying bake attempts that were less successful (ref. OMC elog 379), particularly as the oven seems to have ramped to an overtemperature state. I am a bit mystified, and I would like to see the oven tuning characterized to a greater extent than I have had time and bandwith to complete within this effort.
  • The maximum datalogger temperature was 104 °C, and the duration of the soak (94 °C or higher) was 68 minutes. This was in contrast to a programmed soak of 2.5 hours and a programmed setpoint of 84 °C.
  • The drive thermocouple did appear to be under-reporting temperature relative to the datalogger thermocouple, but this was not confirmed during the soak period. Neither thermocouple was calibrated as part of this effort.


 

Attachment 1: OMC_ABO_PZT_Curing_Bake_effort_201906_thru_201909.xls
Attachment 2: production_cure_bake_pzt_assys_9_and_10_20190927.png
production_cure_bake_pzt_assys_9_and_10_20190927.png
  381   Mon Sep 30 23:16:53 2019 KojiOpticsGeneralOMC(004): PZT sub-assembly gluing (#9/#10)

Friday: [Stephen, Koji]

As the oven setting has qualified, we brought the PZT assys in the air bake oven.

Monday: [Stephen, Shruti, Koji]

We brought the PZT assys to the clean room. There was not bonding between the flexture and the PZT subassy (Good!). Also the bonding o at each side looks completely wetted and looks good. The package was brought to the OMC lab to be tested in the optical setup.

Attachment 1: IMG_8950.jpeg
IMG_8950.jpeg
Attachment 2: IMG_8953.jpeg
IMG_8953.jpeg
Attachment 3: IMG_8954.jpeg
IMG_8954.jpeg
Attachment 4: IMG_8955.jpeg
IMG_8955.jpeg
  380   Thu Sep 26 17:33:52 2019 StephenGeneralGeneralDirty ABO test run prior to PZT Subassembly Bonding - ABO is Ready!

Follow up on OMC elog 379

I was able to obtain the following (dark blue) bake profile, which I believe is adequate for our needs.

The primary change was a remounting of the thermocouple to sandwich it between two stainless steel masses. The thermocouple bead previously was 1) in air and 2) close to the oven skin.

Attachment 1: image_showing_20190924_abo_qualifying_bake.png
image_showing_20190924_abo_qualifying_bake.png
  379   Tue Sep 24 12:19:20 2019 StephenGeneralGeneral Dirty ABO test run prior to PZT Subassembly Bonding

The 40m Bake Lab's Dirty ABO's OMEGA PID controller was borrowed for another oven in the Bake Lab (sound familiar? OMC elog 377), so I have had to play with the tuning and parameters to recover. This bake seemed to inadequately match the intended temperature profile for some reason (intended profile is shown by plotting prior qualifying bake for comparison).

The parameters utilized here are exactly matching the prior qualifying bake, except that the autotuning may have settled on different parameters.

Options to proceed, as I see them, are as follows:

  1. reposition the oven's driving thermocouple closer to the load and attempt to qualify the oven again overnight
  2. retune the controller and attempt to qualify the oven again overnight
  3. proceed with current bake profile, except monitor the soak temperature via data logger thermocouple and intervene if temperature is too high by manually changing the setpoint.

 

Attachment 1: image_showing_20190923_abo_qualifying_bake.png
image_showing_20190923_abo_qualifying_bake.png
  378   Mon Sep 23 21:29:51 2019 KojiOpticsGeneralOMC(004): PZT sub-assembly gluing (#9/#10)

[Stephen, Shruti, Koji]

We worked on the gluing of the PZT sub-assy (#9 and #10) along with the designed arrangement by Shruti (OMC ELOG 374).

The detailed procedures are described in E1300201 Section 6.2 PZT subassembly and Section 7.3 EP30-2 gluing.

We found that the PZTs, which were debonded from the previous PZT sub assy with acetone, has some copper wires oxidized. However, we confirmed that this does not affect the conductivity of the wires, as expected.

The glue test piece cooked in the toaster oven showed excellent curing. GO SIGNAL

Stephen painted the PZT as shown in Attachment 1.

The fixtures were closed with the retaining plate and confirmed that the optics are not moving in the fixtures.


At this point, we checked the situation of the air-bake oven. And we realized that the oven controller was moved to another vacuum oven and in use with a different setting.

Stephen is going to retrieve the controller to the air bake oven and test the temp profile overnight. Once we confirm the setting is correct, the PZT sub assys will be heat cured in the oven.  Hopefully, this will happen tomorrow. Until then, the sub-assys are resting on the south flow bench in the cleanroom.

Attachment 1: IMG_8933.jpg
IMG_8933.jpg
Attachment 2: IMG_8934.jpg
IMG_8934.jpg
  377   Wed Sep 18 23:38:52 2019 StephenGeneralGeneralDirty ABO ready for PZT Subassembly Bonding

The 40m Bake Lab's Dirty ABO's OMEGA PID controller was borrowed for another oven in the Bake Lab, so I have had to play with the tuning and parameters to recover a suitable bake profile. This bake is pictured below (please excuse the default excel formatting).

I have increased the ramp time, temperature offset, and thermal mass within the oven; after retuning and applying the parameters indicated, the rate of heating/cooling never exceeds .5°C/min.

Expected parameters:  Ramp 2.5 hours Setpoint 1 (soak temperature) 94 °C no additional thermal mass
Current parameters:    Ramp 4 hours Setpoint 1 (soak temperature) 84 °C Thermal mass added in the form of SSTL spacers (see photo)


The ABO is controlled by a different temperature readout from the data logger used to collect data; the ABO readout is a small bead in contact with the shelf, while the data logger is a lug sandwiched between two stainless steel masses upon the shelf. I take the data logger profile to be more physically similar to the heating experienced by an optic in a gluing fixture, so I feel happy about the results of the above bake.

I plan to add the data source file to this post at my earliest convenience.

Attachment 1: index.png
index.png
  376   Wed Sep 18 23:16:06 2019 StephenSupplyGeneralItems staged at 40m Bake Lab for PZT Subassembly Bonding

The following items are presently staged at the 40m Bake Lab (see photo indicating current location) (noting items broght by Koji as well):

  1. Bonding fixtures, now modified with larger washers to constrain springs, and with modification from OMC elog 358.
  2. Curved Mirrors and Tombstones as selected by Shruti in OMC elog 374.
  3. PZTs as debonded from first iteration subassemblies (SN 12 and SN 13)
  4. Epoxy-cure-testing toaster oven
  5. Other items I can't think of but will populate later  =D

The following item is in its home in Downs 303 (Modal Lab)

  1. EP30-2 epoxy (expiration 2020 Jan 22) with full kit (tracked in PCS via location update [LINK])

 

Attachment 1: IMG_5216.JPG
IMG_5216.JPG
Attachment 2: IMG_5215.JPG
IMG_5215.JPG
  375   Wed Sep 18 22:30:11 2019 StephenSupplyGeneralEP30-2 Location and Status

Here is a summary of the events of the last week, as they relate to EP30-2.

1) I lost the EP30-2 syringes that had been ordered for the OMC, along with the rest of the kit.

  • Corrective action: Found in the 40m Bake Lab garbing area.
  • Preventative action: log material moves and locations in the OMC elog
  • Preventative action: log EP30-2 moves and locations in PCS via location update [LINK]
  • Preventative action: keep EP30-2 kit on home shelf in Modal Lab unless kit is in use

2) The EP30-2 syringes ordered for the OMC Unit 4 build from January had already expired, without me noticing.

  • Corrective action: Requested LHO ship recently-purchased EP30-2 overnight
  • Preventative action: log expiration dates in OMC elog
  • Preventative action: begin purchasing program supported by logistics, where 1 syringe is maintained on hand and replaced as it expires

3) LHO shipped expired epoxy on Thursday. Package not opened until Monday.

  • Corrective action: Requested LHO ship current EP30-2 overnight, this time with much greater scrutiny (including confirming label indicates not expired)
  • Preventative action: Packages should be opened, inspected, and received in ICS or Techmart on day of receipt whenever possible.

4) Current, unopened syringe of EP30-2 has been received from LHO. Expiration date is 22 Jan 2020. Syringe storage has been improved. Kit has been docked at its home in Downs 303 (Modal Lab) (see attached photo, taken before receipt of new epoxy).

Current Status: Epoxy is ready for PZT + CM subassembly bonding on Monday afternoon 23 September.

Attachment 1: IMG_5217.JPG
IMG_5217.JPG
  374   Thu Sep 5 15:40:42 2019 shrutiOpticsConfigurationPZT Sub-Assembly

Aim: To find the combinations of mounting prism+PZT+curved mirror to build two PZT sub-assemblies that best minimises the total vertical beam deviation.

(In short, attachment 1 shows the two chosen sets of components and the configuration according which they must be bonded to minimize the total vertical angular deviation.)

The specfic components and configuration were chosen as follows, closely following Section 2.3.3 of T1500060:

Available components:

Mounting prisms: 1,2,12,14,15 (Even though there is mention of M17 in the attachments, it can not be used because it was chipped earlier.)

PZTs: 12,13

Curved mirrors: 10,13

 

Method:

For a given choice of prism, PZT and mirror, the PZT can be placed either at 0deg or 180deg, and the mirror can rotated. This allows us to choose an optimal mirror rotation and PZT orientation which minimises the vertical deviation.

Total vertical angle = $\theta_{v, prism} +\theta_{v,wedge} +\theta_{v,mirror}$

\theta_{v, prism} was measured by Koji as described in elog 369.

\theta_{v, wedge} [\text{arcsec}] = \theta_{PZT} \sin{\frac{\pi \phi_{PZT}}{180}},             \theta_{PZT}, \phi_{PZT} are the wedge angle and orientation respectively and were measured earlier and shown in elog 373 . 

\theta_{v, mirror} [\text{arcsec}] = \frac{180 \times 3600 \times d}{\pi R_{RoC}} \times \sin{\frac{\pi (\phi-\phi_{ROT})}{180}},               The measurement of the location of the curvature bottom (d, \phi) of the mirrors is shown in elog 372 . The optimal \phi_{ROT} is to be found.

 

These steps were followed:

  1. For every combination of prism, PZT, and mirror, the total vertical deviation was minimized with respect to the angle of rotation of the curved mirror computationally (SciPy.optimize.minimize). The results of this computation can be found in Attachment 2: where Tables 1.1 and 2.1 show the minimum achievable deviations for mirrors C10 and C13 respectively, and Tables 1.2 and 2.2 show the corresponding angle of rotation of the mirrors \phi_{ROT} .
  2. From the combinations that show low total deviations (highlighted in red in Attachment 2), the tolerances for 5 arcsec and 10 arcsec deviations with mirror rotation were calculated, and is shown in Tables 1.3, 1.4, 2.3, 2.4 of Attachment 2.
  3. While calculating the tolerances, the dependence of the vertical deviations with rotation were also plotted (refer Attachment 3).
  4. Two sets from available components with low total deviation and high tolerance were chosen. 

 

Result:

These are the ones that were chosen:

  1. M14 + PZT13 at 0deg + C13 rotated by 169deg anticlockwise (tot vertical dev ~ -3 arcsec)
  2. M12 + PZT12 at 0deg + C10 rotated by 88deg clockwise (tot vertical dev ~0 arcsec)

The method of attaching them is depicted in Attachment 1.

 

Attachment 1: Diagrams_SubAssembly.pdf
Diagrams_SubAssembly.pdf Diagrams_SubAssembly.pdf
Attachment 2: C10_C13_Combinations.pdf
C10_C13_Combinations.pdf C10_C13_Combinations.pdf
Attachment 3: Plots_Config_Tolerance.pdf
Plots_Config_Tolerance.pdf
  373   Thu Aug 29 11:51:49 2019 shrutiOpticsCharacterizationWedging of the debonded PZTs - Calculation

Using the measurements of PZTs 12,13 taken by Stephen, I estimated the wedging angle and orientation following Section 2.3.1 of T1500060. The results can be found in Attachment1 and is summarised as follows.

For PZT 12, PZT 13 respectively:

Avg. height = 2.0063 mm, 2.0035 mm

Wedge direction (from the same direction as in the doc: positive right) = 120 deg, 120 deg

Wedge angles = 45.8 arcsec, 30.6 arcsec

 

This was done assuming that the measurements were taken uniformly at intervals of 60deg along the inner rim of the PZT. The diameter (2r) of the inner rim, according to T1500060, is 9mm. The measured heights were fitted with the function

h = h_0 + \tan(\Omega)\text{ }r(1-\cos(\theta - \alpha))

as depicted in Attachment2 to find wedging angle (\Omega) and orientation (\alpha).

Quote:

Wedge and thickness measurements of PZTs 12 and 13 took place after debonding and cleaning - results are shown in the first image (handwritten post-it format).

These thickness measurements seem to have come back thinner than previous measurements. It is possible that I have removed some PZT material while mechanically removing glue. It is also possible that there is systematic error between the two sets of measurements. I did not run any calculations of wedge ange or orientation on these data.

Note that cleaning of debonded PZTs involved mechanically separating glue from the planar faces of PZTs. The second image shows the razer blade used to scrape the glue away.

There were thick rings of glue where there had been excess squeezed out of the bond region, and there was also a difficult-to-remove bond layer that was thinner. I observed the presence of the thin layer by its reflectivity. The thick glue came off in patches, while the thin glue came off with a bit of a powdery appearance. It was hard to be certain that all of the thin bond layer came off, but I made many passes on each of the faces of the 2 PZTs that had been in the bonded CM assemblies. I found it was easiest to remove the glue in the bonded

I was anticipating that the expected 75-90 micron bond layer would affect the micrometer thickness measurements if it was still present, but I did not notice any irregularities (and certainly not at the 10 micron level), indicating that the glue was removed successfully (at least to the ~1 micron level).

 

Quote:

Yesterday I measured the thickness of the PZTs in order to get an idea how much the PZTs are wedged.

For each PZT, the thickness at six points along the ring was measured with a micrometer gauge.
The orientation of the PZT was recognized by the wire direction and a black marking to indicate the polarity.

A least square fitting of these six points determines the most likely PZT plane.
Note that the measured numbers are assumed to be the thickness at the inner rim of the ring
as the micrometer can only measure the maximum thickness of a region and the inner rim has the largest effect on the wedge angle.
The inner diameter of the ring is 9mm.



The measurements show all PZTs have thickness variation of 3um maximum.

The estimated wedge angles are distributed from 8 to 26 arcsec. The directions of the wedges seem to be random
(i.e. not associated with the wires)



As wedging of 30 arcsec causes at most ~0.3mm spot shift of the cavity (easy to remember),
the wedging of the PZTs is not critical by itself. Also, this number can be reduced by choosing the PZT orientations
based on the estimated wedge directions --- as long as we can believe the measurements.



Next step is to locate the minima of each curved mirror. Do you have any idea how to measure them?

 

 

Attachment 1: PZT_Wedging_Results.pdf
PZT_Wedging_Results.pdf
Attachment 2: PZT_Wedging_Calc.pdf
PZT_Wedging_Calc.pdf
  372   Fri Aug 23 11:11:44 2019 shrutiOpticsCharacterizationFinding the curvature bottom

I attempted to fit the data taken by Koji of the beam spot precession at the CCD in order to find the location of the curvature bottom in terms of its distance (d) and angle (\phi) from the centre of the mirror. This was done using the method described in a previous similar measurement  and Section 2.1.3 of T1500060.

Initially, I attempted doing a circle_fit on python as seen in Attachment 1, and even though more points seem to coincide with the circle, Koji pointed out that the more appropriate way of doing it would be to fit the following function:

f(i, \theta, r, \phi) = \delta_{i,0} [r \cos(\theta+\phi) + x_c] + \delta_{i,1} [r \sin(\theta+\phi) +y_c]

since that would allow us to measure the angle \phi more accurately; \phi is the anti-clockwise measured angle that the curvature bottom makes with the positive x direction.

As seen on the face of the CCD, x is positive up and y is positive right, thus, plotting it as the reflection (ref. Attachment 2) would make sure that \phi is measured anti-clockwise from the positive x direction.

 

The distance from the curvature bottom is calculated as 

d = \frac{rR}{2L}

r: radius of precession on CCD screen (value obtained from fit parameters, uncertainty in this taken from the std dev provided by fit function)

R: radius of curvature of the mirror 

L: Distance between mirror and CCD

 

R = 2.575 \pm 0.005 m (taken from testing procedure doc referenced earlier) and L = 0.644 \pm 0.005 m (value taken from testing doc, uncertainty from Koji)

  d (mm) \phi (deg)
C7 0.554 \pm 0.004 -80.028 \pm 0.005
C10 0.257 \pm 0.002 -135.55 \pm 0.02
C13 0.161 \pm 0.001 -79.31 \pm 0.06

 

Attachment 1: CircleFit.pdf
CircleFit.pdf
Attachment 2: SineFit.pdf
SineFit.pdf
  371   Thu Aug 22 12:35:53 2019 StephenOpticsCharacterizationWedging of the debonded PZTs 2019 August

Wedge and thickness measurements of PZTs 12 and 13 took place after debonding and cleaning - results are shown in the first image (handwritten post-it format).

These thickness measurements seem to have come back thinner than previous measurements. It is possible that I have removed some PZT material while mechanically removing glue. It is also possible that there is systematic error between the two sets of measurements. I did not run any calculations of wedge ange or orientation on these data.

Note that cleaning of debonded PZTs involved mechanically separating glue from the planar faces of PZTs. The second image shows the razer blade used to scrape the glue away.

There were thick rings of glue where there had been excess squeezed out of the bond region, and there was also a difficult-to-remove bond layer that was thinner. I observed the presence of the thin layer by its reflectivity. The thick glue came off in patches, while the thin glue came off with a bit of a powdery appearance. It was hard to be certain that all of the thin bond layer came off, but I made many passes on each of the faces of the 2 PZTs that had been in the bonded CM assemblies. I found it was easiest to remove the glue in the bonded

I was anticipating that the expected 75-90 micron bond layer would affect the micrometer thickness measurements if it was still present, but I did not notice any irregularities (and certainly not at the 10 micron level), indicating that the glue was removed successfully (at least to the ~1 micron level).

 

Quote:

Yesterday I measured the thickness of the PZTs in order to get an idea how much the PZTs are wedged.

For each PZT, the thickness at six points along the ring was measured with a micrometer gauge.
The orientation of the PZT was recognized by the wire direction and a black marking to indicate the polarity.

A least square fitting of these six points determines the most likely PZT plane.
Note that the measured numbers are assumed to be the thickness at the inner rim of the ring
as the micrometer can only measure the maximum thickness of a region and the inner rim has the largest effect on the wedge angle.
The inner diameter of the ring is 9mm.



The measurements show all PZTs have thickness variation of 3um maximum.

The estimated wedge angles are distributed from 8 to 26 arcsec. The directions of the wedges seem to be random
(i.e. not associated with the wires)



As wedging of 30 arcsec causes at most ~0.3mm spot shift of the cavity (easy to remember),
the wedging of the PZTs is not critical by itself. Also, this number can be reduced by choosing the PZT orientations
based on the estimated wedge directions --- as long as we can believe the measurements.



Next step is to locate the minima of each curved mirror. Do you have any idea how to measure them?

 

Attachment 1: IMG_4775.JPG
IMG_4775.JPG
Attachment 2: IMG_4770.JPG
IMG_4770.JPG
  370   Mon Jul 1 12:49:42 2019 KojiOpticsCharacterizationScattering measurement of A and C mirrors

Liyuan's scattering measurement for the A and C mirrors.

Attachment 1: omc_cm_tis_062419.pdf
omc_cm_tis_062419.pdf omc_cm_tis_062419.pdf
Attachment 2: omc_prism_tis_062419.pdf
omc_prism_tis_062419.pdf omc_prism_tis_062419.pdf omc_prism_tis_062419.pdf
  369   Mon Jul 1 12:38:49 2019 KojiOpticsCharacterizationA and M prisms perpendicularity measurement

[Stephen, Koji]

The perpendicularity of some of the A and M prisms were tested.

Results

- The measurement results are listed as Attachment 1 and 2 together with the comparisons to the measurement in 2013 and the spec provided from the vendor.
- Here, the positive number means that the front side of the prism has larger angle than 90deg for the air side. (i.e. positive number = facing up)
- The RoC of the curved mirrors is 2.5m. Therefore, roughly speaking, 83arcsec corresponds to ~1mm beam spot shift. The requirement is 30 arcsec.
- The A prisms tend to have positive and small angle deviations while the M prisms to have negative and large (~50arcsec) angle deviations.
- The consistency: The measurements in 2013 and 2019 have some descrepancy but not too big. This variation tells us the reliability of the measurements, say +/-30arcsec.

Setup

- The photos of the setup is shown as Attachments 3/4/5. Basically this follows the procedure described in Sec 2.2.2 of T1500060.
- The autocollimator (AC) is held with the V holders + posts.
- The periscope post for the turning Al mirror was brought from Downs by Stephen.
- The turning mirror is a 2" Al mirror. The alignment of the turning mirror was initially aligned using the retroreflection to the AC. Then the pitching of the holder was rotated by 22.5deg so that the AC beam goes down to the prism.
- The prism is held on a Al mirror using the post taken from a prism mount.
- If the maximum illumination (8V) is used, the greenish light becomes visible and the alignment becomes easier.
- There are two reflections 1) The beam which hits the prism first, and then the bottom mirror second, 2) The beam which hits the bottom mirror first and then the prism second. Each beam gains 2 theta compared to the perfect retroreflection case. Therefore the two beams have 4 theta of their relative angle difference. The AC is calibrated to detect 2 theta and tells you theta (1div = 1 arcmin = 60 arcsec). So just read the angle defferencein the AC and divide the number by 2 (not 4).

Attachment 1: A_prism.png
A_prism.png
Attachment 2: M_prism.png
M_prism.png
Attachment 3: P_20190627_222658.jpg
P_20190627_222658.jpg
Attachment 4: setup2.JPG
setup2.JPG
Attachment 5: M01_1_id.JPG
M01_1_id.JPG
Attachment 6: A14_meas.JPG
A14_meas.JPG
  368   Mon Jun 24 12:54:58 2019 KojiCleanGeneralHEPA BOOTH

https://www.airscience.com/purair-flow-laminar-flow-cabinets

  367   Tue May 28 12:14:20 2019 StephenOpticsGeneralCM PZT Assembly Debonding of EP30-2 in Acetone

[LiyuanZ, StephenA]

Downs B119

Summary: Beginning on 20 May 2019, two CM PZT assemblies were soaked in Acetone in an effort to debond the EP30-2 bonds between tombstone-PZT and between PZT-optic. Debonding was straightforward after 8 days of soaking. 24 hours of additional acetone soaking will now be conducted in an attempt to remove remnant EP30-2 from bonding surfaces.

Procedure: The assemblies were allowed to soak in acetone for 8 days, with acetone level below the HR surface of the optic. No agitation of the solution, mechanical abrasion of the bond, or other disturbance was needed for the bond to soften.

GariLynn contributed the glassware and fume hood, and advised on the process (similar to debonding of CM and PZT from OMC SN002 after damaging event). The equipment list was (WIP, more detail / part numbers will be gathered today and tomorrow):

  • crystallizing dish (no spout, like a deep petri dish)
  • curved lid
  • wax sheet (to seal)
  • acetone
  • fume hood

Results: Today, 28 May 2019, I went to the lab to check on the optics after 8 days of soaking. Liyuan had monitored the acetone level during the first 4 days, topping up once on 24 May. All bonds were fully submerged for 8 days.

There were 2 assemblies soaked in one crystallizing dish. Debonded assemblies - ref OMC eLOG 328 for specified orientations and components:

PZT Assy #9 - ref. OMC eLOG 334 - M17+PZT#12+C10

PZT Assy #7 - ref. OMC eLOG 332 - M1+PZT#13+C13

PZT Assy #7 was investigated first.

  • C13 was removed with no force required.
  • PZT#13 was removed with no force required.
  • EP30-2 remained at the bond surfaces and tracing the diameters of each bond on each of the 3 bonding surfaces of the PZT and tombstone - these components were returned to the dish to soak.
  • No EP30-2 remained on the surface of the curved mirror - C13 was removed and stored.

A video of removal of C10 and PZT#12 from PZT Assy #9 was collected (See Attachment 8), showing the ease with which the debonded components could be separated.

  • C10 was removed with no force required.
  • A slight force - applied by gripping the barrel of the PZT and pushing with the index finger on the surface of the tombstone - was required to separate PZT#12 from M17,
    • likely due to excess glue at the barrel of the PZT
  • EP30-2 remained at the bond surfaces and tracing the diameters of each bond on each of the 3 bonding surfaces of the PZT and tombstone - these components were returned to the dish to soak.
  • No EP30-2 remained on the surface of the curved mirror - C13 was removed and stored.

Photos and video have been be added to supplement this report (edit 2019/07/08).

 

Attachment 1: omc367_IMG_3499_omc_removal_c13_from_CM7.JPG
omc367_IMG_3499_omc_removal_c13_from_CM7.JPG
Attachment 2: omc367_IMG_3500_omc_removal_pzt13_from_CM7.JPG
omc367_IMG_3500_omc_removal_pzt13_from_CM7.JPG
Attachment 3: omc367_IMG_3501_omc_removal_pzt13_from_CM7_thickness.JPG
omc367_IMG_3501_omc_removal_pzt13_from_CM7_thickness.JPG
Attachment 4: omc367_IMG_3505_omc_removal_M1_from_CM7.JPG
omc367_IMG_3505_omc_removal_M1_from_CM7.JPG
Attachment 5: omc367_IMG_3507_omc_removal_c10_from_CM9.JPG
omc367_IMG_3507_omc_removal_c10_from_CM9.JPG
Attachment 6: omc367_IMG_3512_omc_removal_pzt12_from_CM9.JPG
omc367_IMG_3512_omc_removal_pzt12_from_CM9.JPG
Attachment 7: omc367_IMG_3515_omc_removal_m17_from_CM9.JPG
omc367_IMG_3515_omc_removal_m17_from_CM9.JPG
Attachment 8: omc367_IMG_3506_omc_removal_of_c10_and_pzt12_from_CM9.MOV
  366   Thu May 23 23:27:38 2019 KojiOpticsCharacterizationIGHQEX3000 high power test

LaserComponents IGHQEX3000 (Cage B2: Serial# B1-23) was exposed to the beam with the optical power from 1.6mW to 332mW.
After each illumination, the dark current and the dark noise level were measured. Also the photo image of the PD surface was taken each time.

- No significant change of the dark current after each illumination.

- No significant change of the dark noise after each illumination.

- No visible change of the surface observed.

(During this dark noise measurement, the current amp gain was set to be 1e8 V/A, instead of 1e7 for the measurements yesterday.)

Attachment 1: HQEPD_high_power_test.pdf
HQEPD_high_power_test.pdf
Attachment 2: pd_images.png
pd_images.png
  365   Thu May 23 01:42:46 2019 KojiOpticsCharacterizationC30665 high power test

An Excelitas C30665 PD with the cap removed (SN07 in Case H slot #2) was exposed to the beam with the optical power of 1.4mW to 334mW.
After each illumination, the dark current and the dark noise level were tested. Also the photo image of the PD surface was taken each time.

- No significant change of the dark current after each illumination.

- No significant change of the dark noise after each illumination.

- No visible change of the surface observed.

 

Attachment 1: C30665_high_power_test.pdf
C30665_high_power_test.pdf
Attachment 2: pd_surface.jpg
pd_surface.jpg
  364   Wed May 22 07:31:37 2019 KojiOpticsConfigurationCamera test (DCPD high power test)

C30665 (3mm) camera test. The camera was Canon PowerShot G7X MkII. Exposure 1/15s, F 5.6, ISO 125, MF (~the closest), no zoom.
This image was taken before the beam illumination. Will tune the green lighting to have some gradient on the surface so that we can see any deformation of the surface.

 

Attachment 1: 20190521201838_IMG_7939_2.jpg
20190521201838_IMG_7939_2.jpg
  363   Mon May 20 19:53:17 2019 KojiOpticsConfigurationDCPD high power test

We want to perform a damage test of OMC DCPDs with high power beam. The OMC DCPD is the 3mm InGaAs photodiodes with high quantum efficiency, delivered by Laser Components.
The sites want to know the allowed input power during the OMC scan for beam mode analysis. The nominal bias voltage of the PDs is +12V. Therefore, 30mA of photocurrent with the transimpedance of 400 Ohm is already enough to saturate the circuit. This means that the test is intended to check the damage of the photodiode mainly by the optical power.

The test procedure is as follows:

1. Illuminate the diode with certain optical power.
2. Measure the dark current and dark noise of the PD with no light on it.
3. Check the condition of the PD surface with a digital camera.
4. Repeat 1~3 with larger optical power.

The beam from an NPRO laser is delivered to the photodiode. The maximum power available is 300~400mW. The beam shape was regulated to have the beam radius of ~500um.

- When the PD is exposed to the high power beam, the circuit setup A) is used. This setup is intended to mimic the bias and transimpedance configuration used in the DCPD amp at the site.

- When the dark noise is measured, the circuit setup B) is used. This setup is low noise enough to measure the dark noise (and current) of the PD.

- The test procedure is going to be tested with an Excelitas 3mm InGaAs PD (C30665), and then tested with the high QE PD.

Attachment 1: BIAS.pdf
BIAS.pdf
Attachment 2: P_20190520_204822.jpg
P_20190520_204822.jpg
  362   Thu May 16 12:41:28 2019 ChubGeneralGeneralfire pillow found on optics table

That is an expanding fire pillow, also known as firebrick.  It is used to create a fire block where holes in fire-rated walls are made and prevents lab fires from spreading rapidly to adjacent labs.  I had to pull cable from B254 to our labs on either side during a rather narrow window of time.  Some of the cable holes are partially blocked, making it difficult to reach the cable to them. The cable is then just guided to the hole from a distance.  With no help, it's not possible to see this material getting shoved out of the hole.  I can assure you that I took great pains not to allow the CYMAC coax to fall into any equipment, or drag against any other cables.   

ELOG V3.1.3-